# Covariant Measures of Non-Markovianity in Curved Spacetime

### v1

@@ Abstract
Standard measures of quantum non-Markovianity are usually defined in terms of dynamical maps on a preferred time foliation and therefore do not extend straightforwardly to curved spacetimes, where no global time coordinate exists and causal structure is primary. We develop a covariant framework for open quantum dynamics along arbitrary timelike worldlines by building multi-time quantum processes (process tensors) from overlapping causal diamonds. For an Unruh–DeWitt detector weakly coupled to a scalar field in a Hadamard state, we define a foliation-independent measure of non-Markovianity as the operational distance between the physical process tensor and the convex set of Markovian (CP-divisible) processes. Numerical benchmarks in (1+1)(1{+}1) dimensions compare inertial motion, uniform acceleration, and static and infalling trajectories in Schwarzschild spacetime. Inertial trajectories are found to be almost Markovian, whereas acceleration and curvature generate pronounced long-range temporal correlations and strong non-Markovian behaviour. In Rindler spacetime, acceleration produces horizon-induced memory tails. In Schwarzschild spacetime, near-horizon field correlations cause both static and freely falling observers to experience enhanced memory, which can remain hidden in single-step diagnostics but becomes evident in multi-time protocols and can even be superactivated by combining different time steps. Our results provide, to our knowledge, the first coordinate-independent, operational quantification of quantum memory in relativistic settings. They identify spacetime curvature, horizons, and acceleration as controllable ingredients that can either degrade or be harnessed as resources in relativistic quantum information tasks, including communication and metrology with accelerated detectors and near black holes.
@@ I Introduction
The study of quantum systems interacting with their environments has long revealed that information is rarely lost in a simple, one-way process. Instead, the environment can retain traces of the system’s past, feeding them back at later times in ways that alter its future. Such phenomena, known as non-Markovian dynamics, mark a departure from the memoryless (Markovian) models that dominate quantum optics and open-system theory [1, 2, 3, 4]. In recent years, understanding and quantifying these memory effects has become central not only to quantum information science but also to fundamental physics, where they appear naturally in settings with strong correlations, structured reservoirs, or relativistic motion [5].
Progress in the study of quantum memory effects has been marked by the development of rigorous measures of non-Markovianity. Among the most prominent are the Breuer–Laine–Piilo (BLP) criterion, which identifies increases in the distinguishability of quantum states as signatures of information backflow [6], and the Rivas–Huelga–Plenio (RHP) criterion, which defines non-Markovianity through the loss of complete-positive divisibility of dynamical maps [7, 8]. Related approaches assess memory effects by monitoring the divisibility of dynamical maps or the evolution of system–ancilla entanglement [7, 9]. Collectively, these tools have deepened our understanding of non-Markovian processes in controlled laboratory settings and provided operational insights relevant to quantum control, precision metrology, and error correction [1, 4].
Yet, when these ideas are transplanted into curved spacetime, new difficulties emerge. The defining strength of traditional measures—their reliance on a preferred time parameterization—becomes a weakness in relativistic contexts. Fixed time slices break diffeomorphism invariance, obscuring the true causal structure [10, 11, 12]. Moreover, pairwise notions of information backflow cannot capture the multi-time correlations that are essential in environments shaped by spacetime curvature, where quantum fields correlate events across entire worldlines rather than isolated instants. For example, a quantum detector traversing a black hole horizon can exhibit coherence revivals driven by spacetime curvature—effects that traditional backflow criteria underestimate or miss entirely [13, 14, 15, 16]. These limitations motivate the search for a covariant framework that treats memory as an intrinsic manifestation of spacetime correlations rather than an artifact of foliation.
In this work, we develop a covariant process-tensor formalism for curved spacetime. In Section II, we begin with algebraic quantum field theory and the Hadamard form of two-point functions to derive memory kernels along affine-parameterized worldlines. We demonstrate how these kernels decompose into causal diamonds. We then introduce measures of non-Markovianity based on diamond-norm distances to the nearest Markovian comb and the correlations found in the associated Choi states in Section III. Section IV presents benchmarks with Unruh–DeWitt detectors in Schwarzschild and accelerated frames which recover flat-space results while revealing horizon-induced superactivation of memory, establishing a foliation-independent foundation for quantifying non-Markovian dynamics in relativistic settings.
@@ II Covariant non-Markovianity via process tensors
In this section we construct the operational object that underlies all of our later diagnostics: a covariant process tensor for an open quantum system coupled to a quantum field in curved spacetime.[17, 18, 19, 20] The idea is to describe, in a fully operational and foliation–free way, how an experimenter can probe memory effects by performing interventions along a worldline and reading out the resulting statistics; see also Refs. [1, 2] for broader discussions of quantum non-Markovianity. Technically, the process tensor is a multi–time quantum channel that maps a sequence of local control operations to outcome probabilities or final states, and can be viewed as a special case of a quantum comb or network.[21, 22, 23, 24] Microscopically, it is determined by the unitary dynamics of the system plus field, and in the weak–coupling regime its non-Markovian structure can be expressed in terms of the field’s two–point function, in close analogy with open-system master-equation approaches.[25, 2] When the environment is in a Hadamard state, this two–point function has a universal short–distance singularity and a regular, state-dependent remainder;[26, 12, 27, 28, 29, 30] the corresponding memory kernels inherit the same split. Finally, by grouping contributions in causal diamonds along the worldline, we obtain a description that is manifestly covariant under coordinate changes and independent of any global time slicing, in the spirit of locally covariant quantum field theory and algebraic locality.[31, 10, 32, 33, 34]
@@ II.1 Setup and operational object
We consider a two–level Unruh–DeWitt (UDW) detector $S$ weakly coupled to a real scalar field $\phi$ on a globally hyperbolic spacetime $(\mathcal{M},g_{\mu\nu})$.[14, 35, 36, 16, 37] The detector follows a timelike worldline $x(\lambda)$, parameterized by an affine parameter $\lambda$ along the trajectory. (In concrete examples we will often choose $\lambda$ to coincide with proper time $\tau$, but keeping the parameter general makes covariance manifest.) The interaction Hamiltonian in the interaction picture for the joint system is taken to be
@@
where $g\ll 1$ is a dimensionless coupling constant, $\sigma_{x}$ is a Pauli operator acting on the detector, and $\chi(\lambda)$ is a smooth, compactly supported switching function that localizes the interaction to a finite portion of the trajectory, as is standard in regularized detector models.[36, 35, 16] The field is prepared in a physically reasonable (Hadamard) state $\rho_{E}$ on the algebra of local observables,[26, 27, 29, 37] and the initial joint state is assumed to be factorized,
@@
The factorized initial condition and small coupling are the standard assumptions of the Born (weak–coupling) expansion in open quantum systems.[25, 2, 1] Both will be used explicitly below when we connect the process tensor to the Wightman function of the field.
@@
with $\lambda_{0}$ and $\lambda_{n}$ the initial and final parameter values of the experiment. At each intermediate point $\lambda_{k}$ ($k=1,\dots,n-1$) the experimenter implements a completely positive trace–preserving (CPTP) map $A_{k}$ on $S$; these maps can represent unitary rotations, noisy gates, measurements followed by feed–forward, or simply the identity channel.[1, 2] The field $E$ is not directly controlled: between intervention points, the combined system evolves unitarily under $H_{\mathrm{int}}(\lambda)$.
The process tensor is the multi–time map that takes this entire sequence of operations as input and returns the statistics of any final measurement, or equivalently the final state of $S$ at $\lambda_{n}$.[18, 17, 19, 20] We write it as
@@
where $\rho_{S}^{(\mathrm{out})}$ is the state of the detector at $\lambda_{n}$ after the specified interventions. More generally, if each $A_{k}$ is an instrument with classical outcomes, $\Upsilon_{n:1}$ assigns joint probabilities to outcome strings.[24] From an operational point of view, $\Upsilon_{n:1}$ is the most general object that encodes how different choices of local control along the worldline lead to different observed statistics; memory effects correspond to the failure of $\rho_{S}^{(\mathrm{out})}$ to depend only on the most recent intervention.[18, 23, 19]
@@ II.2 Choi representation and construction from the field
To connect the abstract map (4) to the underlying field dynamics, it is convenient to use the Choi–Jamiołkowski representation. For a single quantum channel $\mathcal{E}$ acting on a dd–dimensional Hilbert space, the Choi operator is
@@
where $|\Omega\rangle=\sum_{i=1}^{d}|i\rangle\otimes|i\rangle$ is an (unnormalized) maximally entangled vector.[38, 39] The Choi operator encodes all action of $\mathcal{E}$ and is positive with a partial–trace constraint encoding trace preservation. In the multi–time setting one applies the same isomorphism slot by slot, obtaining a positive operator $\Upsilon$ on the tensor product of “input” and “output” Hilbert spaces associated with each time slice. This operator is often depicted as a quantum comb: a sequence of teeth corresponding to time steps, with legs representing the in/out systems and internal links representing memory.[21, 22, 18]
@@
Applying the Choi isomorphism to the multilinear map (8) produces a positive operator $\Upsilon$ with linear constraints that implement causality (no signalling from the future to the past along the comb).[22, 18, 33, 34] This operator is the fundamental object used in Sec. III to define our distance–based measure of non-Markovianity.
@@ II.3 Environment correlations and Hadamard form
At second order in $g$, the entire influence of the environment on the process tensor is encoded in such two–point functions evaluated between pairs of segments along the worldline, in direct analogy with nonlocal memory kernels in quantum Brownian motion and related models.[25, 2, 1]
In curved spacetime, physically admissible states for the field are Hadamard states. These are characterized by the universal local structure of the Wightman function near coincidence.[26, 12, 27, 28, 29] Specifically, for any Hadamard state one can write
@@
where $\sigma(x,x^{\prime})$ is Synge’s world function (half the squared geodesic distance between $x$ and $x^{\prime}$), $u$, $v$ and $w$ are smooth biscalars in a neighborhood of the diagonal, and $i\epsilon$ specifies the boundary condition.[40, 30] The leading coefficient $u(x,x^{\prime})$ is related to the Van Vleck–Morette determinant and encodes the universal light-cone singularity.[40, 30] The biscalar $v(x,x^{\prime})$ admits an expansion
@@
with coefficients determined recursively by local curvature invariants (DeWitt coefficients).[30, 28] Finally, the remainder $w(x,x^{\prime})$ is smooth and state-dependent; it carries global information such as thermal or horizon-induced excitations in Hartle–Hawking or Unruh states.[13, 41, 26, 37]
@@
These kernels appear as coefficients in the Born-expanded maps and, in the Choi picture, as weights multiplying system operators (cf. Sec. III B).[25, 18, 19] Physically, $\eta_{i-j}$ quantifies how strongly an excitation or de-excitation of the detector on segment $\gamma_{j}$ influences the dynamics on $\gamma_{i}$ through the field’s correlations. Diagonal entries $\eta_{i-i}$ encode local dissipation and decoherence, while off-diagonal entries $\eta_{i-j}$ with $i\neq j$ signal genuine temporal memory.[1, 2]
@@
The singular part $\eta^{\mathrm{sing}}_{i-j}$ is completely fixed by local geometry: it depends only on the metric through $\sigma$ and the Hadamard coefficients $u$ and vv.[40, 30, 27] In contrast, $\eta^{\mathrm{reg}}_{i-j}$ carries the state–dependent, physically relevant long–range correlations that distinguish, for example, vacuum from thermal or near-horizon states.[13, 41]

### v2

@@ Abstract
Standard diagnostics of quantum non-Markovianity are commonly formulated in terms of dynamical maps defined on a preferred time foliation. This becomes conceptually and operationally ambiguous in curved spacetime, where no global time coordinate exists and causal structure is primary. We develop a covariant framework for open quantum dynamics along arbitrary timelike worldlines by constructing multi-time processes (process tensors) from overlapping causal diamonds. For an Unruh–DeWitt detector weakly coupled to a scalar field in a Hadamard state, we quantify memory by the operational distance between the resulting process tensor and the convex set of Markovian (CP-divisible) combs, yielding a foliation-independent measure of non-Markovianity. Numerical benchmarks in (1+1)(1{+}1) dimensions compare inertial and uniformly accelerated motion, as well as static and infalling trajectories in Schwarzschild spacetime. Inertial motion is nearly Markovian, whereas acceleration, curvature, and horizons generate long-range temporal correlations and strong multi-time memory, including an activation gap: effects that remain weak in single-step diagnostics become detectable—and can be enhanced—by genuinely multi-time protocols. Our results provide an operational route to quantifying quantum memory in relativistic settings and identify acceleration, curvature, and horizons as controllable ingredients for relativistic quantum-information tasks.
@@ I Introduction
The interface of quantum information and gravitation has sharpened concrete questions about what information is operationally accessible in quantum field theory on curved spacetime, from horizon physics to the structure of entanglement and correlations in relativistic settings [23, 51, 54]. A particularly direct operational viewpoint is provided by local probes, particle detectors or observers coupled to a quantum field, whose statistics encode the field’s correlations along a worldline. Because the field can store information about a probe’s earlier interactions and influence later outcomes, the effective reduced dynamics of such probes are generically non-Markovian, and this memory can matter for tasks such as discrimination, control, and metrology in relativistic environments [5, 31].
In flat spacetime, memory effects can often be analyzed with familiar tools: spectral densities, master equations, and influence-functional techniques provide controlled routes to identifying when a Markov approximation is reliable and when it fails [7]. In curved spacetime, however, importing standard non-Markovianity measures faces a conceptual obstacle that is fundamentally geometric. Many widely used diagnostics compare dynamical maps between time slices and implicitly rely on a preferred time parameterization [6, 43]. In general relativity, where global foliations need not exist and causal structure is primary, such constructions risk conflating coordinate choices with physical memory. In addition, two-time criteria are not designed to capture the full hierarchy of multi-time correlations that naturally arise in quantum fields. This limitation becomes especially sharp near horizons, where redshift and backscattering can generate long-lived “memory tails” and revivals that challenge simple Markovian approximations [33, 25]. Relativistic quantum-information studies such as entanglement harvesting and detector correlations in curved backgrounds and in de Sitter also highlight that curvature and acceleration imprint nontrivially on operational correlations [21, 34].
In many detector-based treatments, whether formulated in master-equation language or through influence functionals, one often invokes a Markov or local-in-time approximation (for example, by coarse-graining to a time-local generator) to obtain a tractable evolution for the detector’s reduced state [7, 5, 28]. Near strong redshift, backscattering, or acceleration-induced long correlations, these approximations can fail in ways that are not reliably diagnosed by two-time criteria alone [33, 25]. The framework developed here provides a covariant, multi-time certificate of when memory is operationally present: it quantifies the distinguishability of the physical multi-time process from the closest CP-divisible (memoryless) comb and exposes correlations that become accessible only to genuinely multi-time protocols [40].
In this work we propose a manifestly covariant route to quantifying memory for worldline-local probes. Our starting point is the process-tensor (quantum-comb) description of multi-time dynamics [40]. For field-coupled probes, this object can be viewed as an operational discretization of the Feynman–Vernon influence functional [19]: it encodes the statistics of arbitrary sequences of interventions on the probe and therefore captures precisely the multi-time correlations that two-time diagnostics may miss. To make this construction covariant, we build the process tensor from the field’s correlation functions organized into overlapping causal diamonds along the worldline, so that the resulting multi-time process depends only on proper time and local causal structure rather than any global coordinate choice.
We focus on the canonical Unruh–DeWitt detector model coupled to a scalar field in a Hadamard state. Within the weak-coupling regime, we define non-Markovianity as an operational distinguishability measure: the distance between the physical multi-time process and the convex set of Markovian processes, understood here as CP-divisible combs. This definition elevates “Markovian vs. non-Markovian” from an observer-dependent slicing of dynamics to a statement about which multi-time statistics can (or cannot) be reproduced by a memoryless process compatible with the same intervention structure.
Physically, the resulting “activation gap” reflects the fact that the field can store trajectory-dependent information in correlations that do not appear in any single reduced map, but become accessible only when one correlates interventions across multiple time steps [40, 36]. Operationally, this is witnessed by multi-time channel discrimination: there exist scenarios where any two-time test has near-random success probability, while a genuinely multi-time tester distinguishes the true process from its best Markovian approximation with a significantly higher bias [10, 12].
We then benchmark how geometry controls memory by comparing inertial and uniformly accelerated motion in Minkowski space with static and infalling trajectories in Schwarzschild spacetime (in (1+1)(1{+}1) dimensions). Our benchmarks are designed to answer three concrete questions. (i) Which geometric ingredients (acceleration, curvature, and horizons) produce long-lived temporal correlations along a worldline in a way that survives operational coarse-graining? (ii) To what extent can two-time diagnostics underestimate memory compared to genuinely multi-time protocols built from the full process tensor? (iii) How do these effects differ between static and infalling observers near a black hole horizon? The remainder of the paper develops the covariant construction needed to pose these questions in a foliation-independent way and evaluates them in controlled (1+1)(1{+}1)-dimensional settings.
@@ II Covariant non-Markovianity via process tensors
In this section we construct the multi-time object that underlies all later diagnostics: a covariant description of the detector’s reduced dynamics along a timelike worldline that does not rely on any global time foliation. Conceptually, the relevant quantity for a worldline-local probe is the Feynman–Vernon influence functional restricted to the probe trajectory: it encodes how field correlations couple different portions of the worldline. The process tensor provides an operational discretization of this influence functional [40, 39, 19, 20]: it maps a sequence of interventions on the probe to observable multi-time statistics, and it is therefore sensitive to memory that two-time diagnostics can miss.
To make this construction manifestly covariant, we discretize the worldline into segments $\gamma_{k}=[\lambda_{k-1},\lambda_{k}]$ and group environmental contributions according to the causal diamonds generated by these segments (Fig. 1). The resulting multi-time process depends only on (i) the causal order along the worldline and (ii) field correlation functions pulled back to the trajectory, and is invariant under coordinate reparameterizations. In the weak-coupling regime for an Unruh–DeWitt detector coupled to a scalar field in a Hadamard state, the influence of the field on the process tensor can be expressed in terms of the Wightman function and associated nonlocal memory kernels. This covariant process tensor will be the input for the distance-based non-Markovianity measure introduced in Sec. III.
Operational specification and covariance.
An experiment in our setting is specified by (i) the spacetime and field state $(\mathcal{M},g_{\mu\nu},\rho_{E})$, (ii) a timelike worldline $x(\lambda)$ as a geometric curve, (iii) a switching profile $\chi(\lambda)$ along the curve, (iv) the slot boundaries $\{\lambda_{k}\}$ (equivalently, the associated chain of causal diamonds), and (v) the class of admissible interventions (in the definition, arbitrary CPTP instruments on the detector, possibly with bounded ancilla). The resulting process tensor and the induced non-Markovianity measure are invariant under passive coordinate changes and orientation-preserving reparameterizations of $\lambda$, as well as under diffeomorphisms that preserve the causal-diamond support of interventions (prop. 1). By contrast, changing $\chi(\lambda)$ or $\{\lambda_{k}\}$ changes the operational task and therefore defines a different process tensor rather than a gauge-equivalent description.
@@ II.1 Setup and operational object
We consider a two–level Unruh–DeWitt (UDW) detector $S$ weakly coupled to a real scalar field $\phi$ on a globally hyperbolic spacetime $(\mathcal{M},g_{\mu\nu})$ [51, 32, 47, 33, 15]. The detector follows a timelike worldline $x(\lambda)$, parameterized by an affine parameter $\lambda$ along the trajectory. (In concrete examples we will often choose $\lambda$ to coincide with proper time $\tau$; however, retaining $\lambda$ in its general form illustrates covariance of the framework) The interaction Hamiltonian in the interaction picture for the joint system is taken to be
@@
where $g\ll 1$ is a dimensionless coupling constant, $\sigma_{x}$ is a Pauli operator acting on the detector, and $\chi(\lambda)$ is a smooth, compactly supported switching function that localizes the interaction to a finite portion of the trajectory, as is standard in regularized detector models [47, 32, 33]. The field is prepared in a physically reasonable (Hadamard) state $\rho_{E}$ on the algebra of local observables [30, 27, 18, 15], and the initial joint state is assumed to be factorized,
@@
The factorized initial condition and small coupling are the standard assumptions of the Born (weak–coupling) expansion in open quantum systems [28, 44, 5]. Both will be used explicitly below when we connect the process tensor to the Wightman function of the field.
@@
with $\lambda_{0}$ and $\lambda_{n}$ the initial and final parameter values of the experiment. At each intermediate point $\lambda_{k}$ ($k=1,\dots,n-1$) the experimenter implements a completely positive trace–preserving (CPTP) map $A_{k}$ on $S$; these maps can represent unitary rotations, noisy gates, measurements followed by feed–forward, or simply the identity channel [5, 44]. The field $E$ is not directly controlled: between intervention points, the combined system evolves unitarily under $H_{\mathrm{int}}(\lambda)$.
The process tensor is the multi–time map that takes this entire sequence of operations as input and returns the statistics of any final measurement, or equivalently the final state of $S$ at $\lambda_{n}$ [39, 40, 2, 20]. We write it as
@@
where $\rho_{S}^{(\mathrm{out})}$ is the state of the detector at $\lambda_{n}$ after the specified interventions. More generally, if each $A_{k}$ is an instrument with classical outcomes, $\Upsilon_{n:1}$ assigns joint probabilities to outcome strings [45]. From an operational point of view, $\Upsilon_{n:1}$ is the most general object that encodes how different choices of local control along the worldline lead to different observed statistics; memory effects correspond to the failure of $\rho_{S}^{(\mathrm{out})}$ to depend only on the most recent intervention [39, 35, 2].
@@ II.2 Choi representation and construction from the field
To connect the abstract map [Eq. (4)] to the underlying field dynamics, it is convenient to use the Choi–Jamiołkowski representation. For a single quantum channel $\mathcal{E}$ acting on a dd–dimensional Hilbert space, the Choi operator is
@@
where $|\Omega\rangle=\sum_{i=1}^{d}|i\rangle\otimes|i\rangle$ is an (unnormalized) maximally entangled vector [13, 29]. The Choi operator encodes all action of $\mathcal{E}$ and is positive with a partial–trace constraint encoding trace preservation. In the multi–time setting one applies the same isomorphism slot by slot, obtaining a positive operator $\Upsilon$ on the tensor product of “input” and “output” Hilbert spaces associated with each time slice. This operator is often depicted as a quantum comb: a sequence of teeth corresponding to time steps, with legs representing the in/out systems and internal links representing memory [10, 12, 39].
@@
Applying the Choi isomorphism to the multilinear map (8) produces a positive operator $\Upsilon$ with linear constraints that implement causality (no signalling from the future to the past along the comb)[12, 39, 14, 37]. This operator is the fundamental object used in Sec. III to define our distance–based measure of non-Markovianity.
@@ II.3 Environment correlations and Hadamard form
At second order in $g$, the entire influence of the environment on the process tensor is encoded in such two–point functions evaluated between pairs of segments along the worldline, in direct analogy with nonlocal memory kernels in quantum Brownian motion and related models [28, 44, 5].
In curved spacetime, physically admissible states for the field are Hadamard states. These are characterized by the universal local structure of the Wightman function near coincidence [30, 42, 27, 46, 18]. Specifically, for any Hadamard state one can write
@@
where $\sigma(x,x^{\prime})$ is Synge’s world function (half the squared geodesic distance between $x$ and $x^{\prime}$), $u$, $v$ and $w$ are smooth biscalars in a neighborhood of the diagonal, and $i\epsilon$ specifies the boundary condition [38, 16]. (To avoid confusion, we reserve $\sigma(x,x^{\prime})$ exclusively for Synge’s world function; the width of the switching functions used in the numerical benchmarks will be denoted by $w$ rather than $\sigma$.) The leading coefficient $u(x,x^{\prime})$ is related to the Van Vleck–Morette determinant and encodes the universal light-cone singularity [38, 16]. The biscalar $v(x,x^{\prime})$ admits an expansion
@@
with coefficients determined recursively by local curvature invariants (DeWitt coefficients) [16, 46]. Finally, the remainder $w(x,x^{\prime})$ is smooth and state-dependent; it carries global information such as thermal or horizon-induced excitations in Hartle–Hawking or Unruh states [23, 22, 30, 15].
@@
These kernels appear as coefficients in the Born-expanded maps and, in the Choi picture, as weights multiplying system operators (cf. Sec. III B).[28, 39, 2] Physically, $\eta_{i-j}$ quantifies how strongly an excitation or de-excitation of the detector on segment $\gamma_{j}$ influences the dynamics on $\gamma_{i}$ through the field’s correlations. Diagonal entries $\eta_{i-i}$ encode local dissipation and decoherence, while off-diagonal entries $\eta_{i-j}$ with $i\neq j$ signal genuine temporal memory [5, 44].
@@
The singular part $\eta^{\mathrm{sing}}_{i-j}$ is completely fixed by local geometry: it depends only on the metric through $\sigma$ and the Hadamard coefficients $u$ and $v$ [38, 16, 27]. In contrast, $\eta^{\mathrm{reg}}_{i-j}$ carries the state–dependent, physically relevant long–range correlations that distinguish, for example, vacuum from thermal or near-horizon states [23, 22].